# Environment-Specific People

### v1

@@ Abstract
Despite significant progress in generative image synthesis and full-body generation in particular, state-of-the-art methods are either context-independent, overly reliant to text prompts, or bound to the curated training datasets, such as fashion images with monotonous backgrounds. Here, our goal is to generate people in clothing that is semantically appropriate for a given scene. To this end, we present ESP, a novel method for context-aware full-body generation, that enables photo-realistic inpainting of people into existing “in-the-wild” photographs. ESP is conditioned on a 2D pose and contextual cues that are extracted from the environment photograph and integrated into the generation process. Our models are trained on a dataset containing a set of in-the-wild photographs of people covering a wide range of different environments. The method is analyzed quantitatively and qualitatively, and we show that ESP outperforms state-of-the-art on the task of contextual full-body generation.
Figure 1: ESP generates people in semantically appropriate, environment-inspired clothing: given a 2D pose (row 1) and an environment (row 2); people are inpainted into the photo, based on a context-aware generative model that predicts contextual human parsing maps.
@@ 1 Introduction
In recent years, great progress has been made on all aspects of image generation and editing, e.g., GAN-based methods [53, 56, 49, 28, 54, 55] allow for photo-realistic image synthesis on well-crafted datasets, and diffusion-based models [41, 52, 32] make it possible for the original image content to be seamlessly removed or replaced via inpainting. Due to the rapid developments in generative AI, distinguishing between AI-generated content and real content by analyzing the quality of the generated images alone is getting more difficult each day. In this paper, we focus on full-body generation and inpainting of environment-specific people wearing clothing that is appropriate for a given environment. For example, if the context of the depicted scene is a football field, our method would generate a person dressed in sportswear such as shorts, a T-shirt or a polo shirt, football boots, etc. However, if the context switches to snowy mountains, the inserted person would be wearing a winter jacket, ice skates, or ski attire.
@@
One line of work in full-body generation is centered around using StyleGAN [11, 12] to generate full-body samples. Such methods are usually trained on high-quality data, such as DeepFashion [31], that is curated and hand annotated. To reduce its high dimensionality, data is first normalized and filtered. This is a crucial step that is contributing to the success of StyleGAN-based methods. While the images that are generated in this way are of high-quality, the main issue that we tackle in this paper is the semantic asyncrony between the clothing and the corresponding environment that exists in these datasets due to the unnatural data acquisition settings. I.e., in fashion photography, images are usually captured in a studio setting, where the subject is fully lit and the background is very simple, e.g., a white screen. Therefore there is minimal environmental context available. Also, in case of an outdoor photo-shoot, the link between the environment and the clothing may be atypical and it does not represent what is normally worn on the location well.
@@
The other, more recent line of work, are diffusion-based methods or, typically, large pre-trained latent diffusion models such as Stable Diffusion [41] and its conditional variations (e.g. ControlNet [52]). The main advantage of such methods is the existence of large pre-trained image priors that are usually fine-tuned on human data including full-bodies in the wild. However, in order to generate sufficiently varied samples, these txt2img models require accurate textual descriptions of the desired output. While providing such descriptions may lead to satisfactory results, in this work we assume that there are no textual descriptions of the appropriate clothing available. And instead, we focus on visual information alone, investigating whether the image context can be used as a sole driving source for the generation of environment-specific people.
@@
To this end, we take advantage of both StyleGAN-based and diffusion-based methods, combining their strengths to devise a purely image-based system for our task. As part of our pipeline, contextual cues are first extracted from a given environment photo with a VAE and then fed into a StyleGAN-based human parsing map (HPM) generator that predicts context-aware clothing semantics. HPM training data is filtered and normalized to include only simple poses, making it suitable for training a StyleGAN model. Next, a HPM translation module is fine-tuned using the same dataset and in turn utilized to provide additional guidance to the inpainting process of a large pre-trained Stable Diffusion model, for seamless inpainting of photo-realistic and environment-specific people into the original photographs.
Our method is evaluated quantitatively and qualitatively, with a focus on contextual appropriateness. The results show that our method generates environment-specific people that match the context of the input photograph well, considering the clothing semantics (see Figure 1). Furthermore, our approach outperforms state of the art on the task of contextual full-body generation.
Our contributions are as follows:
@@
  • a dynamic pix2pix-based extension of the computational graph that is fine-tuned in parallel with the StyleGAN generator which enables pose-guided imgage-to-image translation;
  • a system that leverages a pre-trained text-to-image latent diffusion model (Stable Diffusion), allowing for super-resolution imaging and photorealistic inpainting of environment-specific people into existing photographs.
Figure 2: System Overview: (I) an input context image is encoded into the latent space of a VAE, giving a contextual style vector; (II) the latter is then fed, alongside a random vector, into a pose-conditioned StyleGAN-II HPM generator; (III) an HPM translation module is used to achieve fine-grained control over the generated clothing during the inpainting process (using the generic Stable Diffusion LDM).
@@ 2 Related work
From their advent, Generative Adversarial Networks (GAN) [13] have proven to be a very effective tool for synthesizing realistic images. StyleGAN-based methods [20, 21, 23, 4] have been used to generate photo-realistic imagery of a variety of objects and scenes [50, 51] , including humans [26, 45]. More recently, large latent diffusion models have been in the spotlight [37] due to their unparalled image generation capabilities, conditioned on text (txt2img)  [41, 38, 39, 1, 42] and additional input controls  [52].
@@
StyleGAN-based state-of-the-art for unconditional full-body generation include InsetGAN [11] and StyleGAN-Human [12]. Both methods use well prepared datasets that are either proprietary or contain fashion imagery. InsetGAN composes multiple GANs to handle the complex task of generating high resolution full body human images. One GAN is trained to generate the global structure, i.e., the body, while the remaining GANs are each trained separately for different local regions, e.g., face or shoes. The latent spaces of different GANs are jointly optimized for a seamless combination to synthesize the whole person. StyleGAN-Human involves training StyleGAN2 [22] on a large, self-curated dataset, consisting of around 230K full body human images. The background in this data is often white or plain, or the link between the background and the clothing is atypical. We use a small dataset of in-the-wild imagery [57] for training, where people are clothed synchronously with the environment, e.g. football gear is worn on a football field and warm winter attire is worn in the snow. Pose and/or appearance information are used to control the process of conditional full-body generation [3, 33, 34, 47, 44]. In a two step process, Lewis et al. [26] first train a dense-pose conditioned StyleGAN2 model with a clothing segmentation branch, after which a new image is synthesized using layered latent space interpolation to transfer a garment to a target body. Grigorev et al. [14] combine 3D parametric mesh-based models of human bodies [36] with neural textures generated by a modified StyleGAN2 model. Ren et al. [40] propose a deep architecture, conditioned on human pose, to combine flow-based operations with an attention mechanism. This architecture has also been extended for use in an unsupervised setting [43]. Sarkar et al. [45] warp encoded part-based latent appearance vectors in a normalized pose-independent space to generated synthetic people in specific poses. Our twist considering this line of work is to tackle the problem from a novel perspective, by incorporating image context as a conditioning through the mapping network. In this way, we generate environemnt-specific people that may be inpainted into the existing photographs, while respecting the scene context in terms of the clothing semantics.
@@
The task of generating new content in specified regions of an existing image is referred to as inpainting. This is a well-studied problem that is particularly suitable for our task [49, 28, 55]. Motivated by the StyleGAN2 modulated convolutions, Zhao et al. [54] improve inpainting reconstruction by using a comodulation layer. Approaches that build on a variational auto-encoder based architecture have also been proposed [56, 53]. Lugmayr et al. [32] improved on the previous methods by harnessing the expressiveness of a pre-trained denoising diffusion probabilistic model (DDPM) [17], which had previously shown impressive results in guided image synthesis in pixel space [9, 35]. Using a pre-trained DDPM as a generative prior, they alter the reverse diffusion iterations by sampling the unmasked regions using the given image information. However, this approach slows down the image generation process of a DDPM. More recently, similar ideas in the realm of latent diffusion models have been explored using denoising diffusion implicit models (DDIM) [48] where the denoising steps are skipped for faster inference. Most recent image generators that have the inpainting capacity use this principle and we take advantage of these models  [41].
@@
Brooks et al. [5] generate both the scene and a pose-compatible human that matches the scene context at 128x128, where the context embeddings are calculated from a 2D pose vector and a noise vector. In their work, the focus is on generating both compatible scenes and humans inside, conditioned on 2D pose embeddings. In contrast, our focus is on inpainting environment-specific people into existing scenes and we calculate context embeddings from the scene context directly. Kulal et al. [25] filter a large-scale video dataset (2.4 million clips) at 256x256 and use it to train a multi-faceted LDM, generating pose-compatible people conditioned on a photo of a reference person and a masked scene. The primary focus of this work is on synthesizing plausible poses given a scene context. For this purpose, two random frames that contain the same human from each video are used during training. At test time, the reference person conditioning may be removed, allowing for the synthesis of scene-compatible humans, given scene information. In contrast, we condition the generation process on a full-body 2D pose rendering (similar to  [8]) and investigate the connection between the scene and its impact on clothing alone, given a fixed 2D pose. Additionally, we only use a small dataset containing approx. 23K images at 200x200, and our results are generated at 512x512. Image quality of our method is on par with state-of-the-art image generators including ControlNet  [52] and Stable Diffusion [41] that many other recent high-quality full-body generation methods build on top of  [30].
@@ 3 Method
Our system is shown in Figure 2. First, context embeddings are extracted from image $\mathcal{I}$ that contains the environmental context $\mathcal{C}$. Then, using the extracted embeddings, small human parsing maps $\mathcal{M}$ are generated conditioned on a 2D skeleton rendering $\mathcal{S}$ (32x64). Next, generated HPMs are then used, with the help of our HPM translation module, as an additional guiding signal for the seamless inpainting of environment-specific people into the existing photographs. In this final step, we take advantage of a large pre-trained LDM (Stable Diffusion for inpainting), allowing for the generation of images at very high resolution (512x512).
@@ 3.1 Context embeddings
We must first find a representation of $\mathcal{I}$ that can be used as the context conditioning to drive the generation of the clothing semantics in the following steps. During training, we have access to the context image containing the human $I_{c}^{h}$ and a corresponding mask of the human $M_{h}$. We dilate the mask with morphological operations and invert it to get the soft background mask $M_{b}$. To remove the foreground human, we apply a pixel-wise mask on the input image, producing the masked image $\mathcal{I}=I_{c}^{h}\odot M_{b}$ Variational auto-encoders (VAE) [24] are good at representing the overall global information present in an image [15], so we choose the compact and normalized latent space of a VAE to reflect the global structure, illumination and color tone of the context image $\mathcal{I}$. We train the VAE with $i_{c}$ where $i_{c}$ is a down-sampled $64\times 64$ version of $\mathcal{I}$ with the loss:
@@
where $\tilde{i}_{c}$ is the reconstructed image from the VAE; $\lambda$’s are the corresponding weights for the different losses; $D_{KL}$ is the KL divergence loss [24]; $\mathcal{N}(x,y)$ is the normal distribution with mean $x$ and variance $y$; and $\mu_{c}$, $\sigma^{2}_{c}$ are the mean and variance of the context image. Once trained, the latent vector $\mathcal{C}$ sampled from the VAE for corresponding context image $\mathcal{I}$ is used for conditioning in the following steps. Note that the reconstruction loss term is applied only to the regions where humans are present and as such we do not require $\mathcal{I}$ to be computed from $I_{c}^{h}$ at inference time. Additionally, $\tilde{i}_{c}$ is subsequently passed through the VAE to remove the small artifacts that may appear in the regions that people typically occupy in our training data.
@@ 3.2 Contextualized generation of HPMs
The HPM generator is additionally conditioned with $\mathcal{S}$ in the image space to enforce the generation of full-body humans in the given pose. This helps us isolate the pose later on in the experiments and focus on exploring the impact that the environment by itself has on the clothing semantics. Motivated by Isola et al. [19], to inject the pose conditioning into the generator, we concatenate the input of different style blocks of the generator with the output of the corresponding layers of a pose encoder. The input to the encoder is an RGB image $\mathcal{S}$ and its layers are designed so that the number of the feature maps and their dimensions match the StyleGAN2 generator (32x64). The generator and the pose encoder are trained end-to-end.
The HPM generator jointly outputs a HPM $\tilde{\mathcal{M}}$ (3 channels) and a blurry background image (3 channels). Note that the blurry background reconstruction from the VAE that is used as the GT image for the generator is rescaled from 64x64 to 32x64 to match the HPM image dimensions. Similar to [11], we observe that GANs do not handle large areas of uniform color well and therefore we use the tall HPM images where the number of the background pixels has been reduced, rather than square images with many background pixels. We are able to do so because our dataset is filtered and it only contains the poses where the width of the person is equal to at most one half of their height.
@@
$\begin{split}L_{S}^{G}=L_{adv}^{D_{\mathcal{M}}}+L_{adv}^{D_{B}}+L_{adv}^{D_{MB}}\\ +\lambda_{3}L_{1}+\lambda_{4}L_{R}+\lambda_{5}L_{perc}\end{split}$ (2)
where $L_{adv}^{D_{\mathcal{M}}}$ and $L_{adv}^{D_{B}}$ represent the adversarial losses that come from the HPM and the blurry background image discriminators respectively; $L_{adv}^{D_{MB}}$ is the loss from an additional discriminator that evaluates the composite of the blurry background and the HPM; $L_{1}$ is the $l_{1}$ loss applied between the generated blurry background and the corresponding generated background $\tilde{i}_{c}$ from the VAE; $L_{perc}$ is a standard perceptual loss commonly used in the literature for generating more detailed images; $L_{R}$ represents the path regularisation [22] that is applied to the generator; $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are the weighting factors for the specific losses.

### v2

# Synthesizing Environment-Specific People in Photographs

@@ Abstract
We present ESP, a novel method for context-aware full-body generation, that enables photo-realistic synthesis and inpainting of people wearing clothing that is semantically appropriate for the scene depicted in an input photograph. ESP is conditioned on a 2D pose and contextual cues that are extracted from the photograph of the scene and integrated into the generation process, where the clothing is modeled explicitly with human parsing masks (HPM). Generated HPMs are used as tight guiding masks for inpainting, such that no changes are made to the original background. Our models are trained on a dataset containing a set of in-the-wild photographs of people covering a wide range of different environments. The method is analyzed quantitatively and qualitatively, and we show that ESP outperforms the state-of-the-art on the task of contextual full-body generation. https://esp.is.tue.mpg.de/
Keywords:
Generative modeling Full-body synthesis Environments
Figure 1: ESP generates people in semantically appropriate, environment-inspired clothing. Given a 2D pose and an environment (top), people are inpainted into the photo based on a context-aware generative model (sample 1, 2, 3).
@@ 1 Introduction
In recent years, great progress has been made on all aspects of image generation and editing. GAN-based methods [57, 60, 52, 30, 58, 59] allow for photo-realistic image synthesis on well-crafted datasets, and diffusion-based models [43, 55, 34] make it possible for the original image content to be seamlessly removed or replaced via inpainting. Rapid advancements in generative AI have resulted in tools that take text prompts and produce photo-realistic images of humans [40, 51]. Constraining the generation requires crafting text prompts to achieve the desired result. In this paper, our goal is to take an image of a scene and a 2D pose and generate a realistic person, in the desired pose, wearing scene-appropriate clothing; no text prompts are required. For example, if the context of the depicted scene is a football field, our method generates people dressed in appropriate athletic wear such as shorts, a jersey, football helmet, etc.; see Figure 1. However, if the context switches to snowy mountains, the inserted person will be wearing a winter jacket, ice skates, or ski attire.
@@
One line of work in full-body generation is centered around using StyleGAN [12, 13] to generate full-body samples. Such methods are trained on high-quality data, such as DeepFashion [33], that is curated and hand annotated. To reduce its high dimensionality, data is first normalized and filtered. This is a crucial step that is contributing to the success of StyleGAN-based methods. While the images that are generated in this way are of high-quality, the generated clothing is typically not correlated with the scene due to the unnatural settings present in the training data. For example, in fashion photography, images are usually captured in a studio setting, where the subject is fully lit and the background is very simple, e.g., a white screen. Therefore, there is minimal environmental context available.
@@
The more recent line of work focuses on diffusion-based methods that generate images from text, e.g. large pre-trained latent diffusion models (LDM) such as Stable Diffusion [43] and its conditional variations (e.g. ControlNet [55]). The main advantage of such methods is the existence of large pre-trained image priors that are usually fine-tuned on human data including full-bodies in the wild. However, in order to generate sufficiently varied samples, these text-to-image (T2M) models require accurate textual descriptions of the desired output. In contrast, here we seek an automated solution and do not assume that textual descriptions of the appropriate clothing are available. Instead, we focus on visual information alone, investigating whether the image context can be used as a sole driving source for the generation of environment-specific people.
@@
To this end, we take advantage of both StyleGAN-based and diffusion-based methods, combining their strengths to devise a purely image-based system for our task. As part of our pipeline, contextual cues are first extracted from a given environment photo with a VAE and, then, fed into a StyleGAN-based human parsing map (HPM) generator that predicts both clothing semantics and the original environment. The key insight is that, when combined with a random vector, the extracted contextual cues form our contextual-style vectors that allow for a stochastic generation of a large variety of context-appropriate clothing. These vectors are connected back to the original environment, as they jointly drive the generation of both the HPMs and the corresponding environment, making the HPM generation context-aware. To disentangle the pose from the context, the generator is further conditioned on 2D pose renderings and the network’s computational graph is extended through an image-to-image (I2I) translation module (similar to pix2pix [20]). Finally, leveraging a large pre-trained T2M LDM for its state-of-the-art image synthesis and inpainting capacities [43], generated HPMs are used as input conditioning and guidance during the inpainting and super-resolution process.
Our method is evaluated quantitatively and qualitatively, demonstrating that our approach outperforms the state of the art on the task of contextual full-body generation based on the following contributions:
  • $\circ$
@@
  • $\circ$
    An extension of the computational graph that is trained end-to-end with the StyleGAN generator which enables pose-guided I2I translation;
  • $\circ$
    A system that leverages a large pre-trained T2M latent diffusion model (Stable Diffusion), allowing for super-resolution imaging and inpainting of the environment-specific people into the existing photographs.
@@ 2 Related work
Generative Adversarial Networks (GAN) [14] have proven to be a very effective tool for synthesizing realistic images. StyleGAN-based methods [24, 22, 23, 5] have been used to generate photo-realistic imagery of a variety of objects and scenes [53, 54], including humans [28, 47]. More recently, large latent diffusion models have been in the spotlight [39] due to their unparalled image generation capabilities, conditioned on text (T2M)  [43, 41, 40, 2, 44] and other controls  [55].
@@
StyleGAN-based state-of-the-art methods for unconditional full-body generation include InsetGAN [12] and StyleGAN-Human [13]. Both methods use well-prepared datasets that are either proprietary or contain fashion imagery. InsetGAN composes multiple GANs to handle the complex task of generating high resolution full body human images. One GAN is trained to generate the global structure, i.e., the body, while the remaining GANs are each trained separately for different local regions, e.g., face or shoes. The latent spaces of different GANs are jointly optimized for a seamless combination to synthesize the whole person. StyleGAN-Human involves training StyleGAN2 [25] on a large, self-curated dataset, consisting of around 230K full body human images. The background in this data is often white or plain, or the link between the background and the clothing is atypical. We use a small dataset of in-the-wild imagery [61] for training, where people are clothed according to the environment, e.g., football gear is worn on a football field and warm winter attire in the snow. Pose and/or appearance information are used to control the process of conditional full-body generation [4, 35, 36, 49, 46]. In a two-step process, Lewis et al. [28] first train a dense-pose conditioned StyleGAN2 model with a clothing segmentation branch, after which a new image is synthesized using layered latent space interpolation to transfer a garment to a target body. Grigorev et al. [15] combine 3D parametric mesh-based models of human bodies [38] with neural textures generated by a modified StyleGAN2 model. Ren et al. [42] propose a deep architecture, conditioned on human pose, to combine flow-based operations with an attention mechanism. This architecture has also been extended for use in an unsupervised setting [45]. Sarkar et al. [47] warp encoded part-based latent appearance vectors in a normalized pose-independent space to generated synthetic people in specific poses. Our twist considering this line of work is to tackle the problem from a novel perspective, by incorporating image context as a conditioning through the mapping network. In this way, we generate environment-specific people that may be inpainted into the existing photographs, while respecting the scene context in terms of the clothing semantics.
@@
The task of generating new content in specified regions of an existing image is referred to as inpainting. This is a well-studied problem that is particularly suitable for our task [52, 30, 59]. Motivated by the StyleGAN2 modulated convolutions, Zhao et al. [58] improve inpainting reconstruction by using a comodulation layer. Approaches that build on a variational auto-encoder based architecture have also been proposed [60, 57]. Lugmayr et al. [34] improved on the previous methods by harnessing the expressiveness of a pre-trained denoising diffusion probabilistic model (DDPM) [18], which had previously shown impressive results in guided image synthesis in pixel space [10, 37]. Using a pre-trained DDPM as a generative prior, they alter the reverse diffusion iterations by sampling the unmasked regions using the given image information. However, this approach slows down the image generation process of a DDPM. More recently, similar ideas in the realm of latent diffusion models have been explored using denoising diffusion implicit models (DDIM) [50], where the denoising steps are skipped for faster inference. Most recent image generators that have the inpainting capacity use this principle and we take advantage of these models  [43].
@@
Brooks et al. [6] generate both the scene and a pose-compatible human that matches the scene context at $128\times 128$, where the context embeddings are calculated from a 2D pose vector and a noise vector. In their work, the focus is on generating both compatible scenes and humans inside, conditioned on 2D pose embeddings. In contrast, our focus is on inpainting environment-specific people into existing scenes and we calculate context embeddings from the scene context directly. Kulal et al. [27] filter a large-scale video dataset (2.4 million clips) at $256\times 256$ and use it to train a multi-faceted LDM, generating pose-compatible people conditioned on a photo of a reference person and a masked scene. The primary focus of this work is on synthesizing plausible poses given a scene context. For this purpose, two random frames that contain the same human from each video are used during training. At test time, the reference person conditioning may be removed, allowing for the synthesis of scene-compatible humans, given scene information. In contrast, we condition the generation process on a full-body 2D pose rendering (similar to  [9]) and investigate the connection between the scene and its impact on clothing alone, given a fixed 2D pose. Additionally, we only use a small dataset containing approx. 23K images at $200\times 200$, and our results are generated at $512\times 512$. Image quality of our method is on par with state-of-the-art image generators including ControlNet [55] and Stable Diffusion [43] that many other recent high-quality full-body generation methods build on top of  [32].
Figure 2: System overview: (I) an input context image is encoded into the latent space of a VAE, giving context embeddings; (II) the latter is then fed, alongside a random vector, creating a contextual style vector, into a pose-conditioned StyleGAN-II HPM generator; (III) generated HPMs are used as input for pretrained Stable Diffusion/CN modules to achieve fine-grained control over the generated clothing during inpainting.
@@ 3 Method
Our method is shown in Figure 2. First, context embeddings are extracted from image $\mathcal{I}$ that contains the environmental context $\mathcal{C}$. Then, using the extracted embeddings, human parsing maps (HPMs) $\mathcal{M}$ are generated conditioned on a 2D skeleton rendering $\mathcal{S}$ ($32\times 64$). The generated HPMs are then used, with the help of our HPM translation module, as an additional guiding signal for inpainting of environment-specific people into the existing photographs. In this final step, we take advantage of a large pre-trained LDM (Stable Diffusion for inpainting), allowing for the generation of images at high resolution ($512\times 512$).
@@ 3.1 Context embeddings
We must first find a representation of $\mathcal{I}$ that can be used as the context conditioning to drive the generation of the clothing semantics in the following steps. During training, we have access to the context image containing the human $I_{c}^{h}$ and a corresponding mask of the human $M_{h}$. We dilate the mask with morphological operations and invert it to get the soft background mask $M_{b}$. To remove the foreground human, we apply a pixel-wise mask on the input image, producing the masked image $\mathcal{I}=I_{c}^{h}\odot M_{b}$
We choose the compact and normalized latent space of a Variational Auto-Encoder (VAE) [26, 16] to reflect the global structure, illumination and color tone of the context image $\mathcal{I}$. We train the VAE with $i_{c}$, where $i_{c}$ is a down-sampled $64\times 64$ version of $\mathcal{I}$ with the loss:
@@
where $\tilde{i}_{c}$ is the reconstructed image from the VAE; $\lambda$s are the corresponding weights for the different losses; $D_{KL}$ is the KL divergence loss [26]; $\mathcal{N}(x,y)$ is the normal distribution with mean $x$ and variance $y$; and $\mu_{c}$, $\sigma^{2}_{c}$ are the mean and variance of the context image. Once trained, the latent vector $\mathcal{C}$ sampled from the VAE for the corresponding context image $\mathcal{I}$ is used for conditioning in the following steps. Note that the reconstruction loss term is applied only to the regions where humans are present, and, as such, we do not require $\mathcal{I}$ to be computed from $I_{c}^{h}$ at inference time.
@@ 3.2 Contextualized generation of HPMs
The HPM generator is additionally conditioned with $\mathcal{S}$ in the image space to enforce the generation of full-body humans in the given pose. This helps us isolate the pose later on in the experiments and focus on exploring the impact that the environment by itself has on the clothing semantics. Motivated by Isola et al. [20], to inject the pose conditioning into the generator, we concatenate the input of different style blocks of the generator with the output of the corresponding layers of a pose encoder. The input to the encoder is an RGB image $\mathcal{S}$ and its layers are designed so that the number of the feature maps and their dimensions match the StyleGAN2 generator ($32\times 64$). The generator and the pose encoder are trained end-to-end. The HPM generator jointly outputs a HPM $\tilde{\mathcal{M}}$ (3 channels) and a blurry background image ($3$ channels). Note that the blurry background reconstruction from the VAE that is used as the GT image for the generator is rescaled from $64\times 64$ to $32\times 64$ to match the HPM image dimensions. Similar to [12], we observe that GANs do not handle large areas of uniform color well and, therefore, we use the tall HPM images, where the number of the background pixels has been reduced, rather than square images with many background pixels. We are able to do so because our dataset is filtered and it only contains the poses where the width of the person is equal to at most one half of their height.
@@
$\begin{split}L_{S}^{G}=L_{adv}^{D_{\mathcal{M}}}+L_{adv}^{D_{B}}+L_{adv}^{D_{MB}}+\lambda_{3}L_{1}+\lambda_{4}L_{R}+\lambda_{5}L_{perc},\end{split}$ (2)
where $L_{adv}^{D_{\mathcal{M}}}$ and $L_{adv}^{D_{B}}$ represent the adversarial losses that come from the HPM and the blurry background image discriminators respectively; $L_{adv}^{D_{MB}}$ is the loss from an additional discriminator that evaluates the composite of the blurry background and the HPM; $L_{1}$ is the $l_{1}$ loss applied between the generated blurry background and the corresponding generated background $\tilde{i}_{c}$ from the VAE; $L_{perc}$ is a standard perceptual loss commonly used in the literature for generating more detailed images [56]; $L_{R}$ represents the path regularisation [25] that is applied to the generator; $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are the weighting factors for the specific losses.